OpenStudy (anonymous):

How can I make (x+2)^2 +y^2=9 into the standard form of the equation of a circle? I know that I have to subtract 0 from y but that leaves (x+2)^2 x+2 when it is supposed to be x-2. Please help )-:

OpenStudy (asnaseer):

use the fact that:\[-(-2)=2\]

OpenStudy (asnaseer):

so:\[(x+2)=(x-(?))\]

OpenStudy (anonymous):

I did that, I was not sure if it was alright for x to be negative.

OpenStudy (asnaseer):

you mean in \((x-a)^2\) you weren't sure if 'a' could be negative?

OpenStudy (anonymous):

the equation was \[(x+2)^{2} +y ^{2} = 9\] I subtracted 0 from y, making the equation \[(x+2)^{2} +(y-0)^{2}=9\] I then multiplied \[(x+2)^{2}\] by -1 which equaled \[(-x-2)^{2}\] because -x is not apart of the formula, I was not sure if that was okay or not

OpenStudy (asnaseer):

no - your last step is incorrect. you need to express (x+2) as (x-?)

OpenStudy (asnaseer):

so:\[x+2=x-(-?)\]what would ? be for this to be true?

OpenStudy (anonymous):

(x-(-2)?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

so:\[(x+2)^2=(x-(-2))^2\]

OpenStudy (anonymous):

okie dokie. I was unaware that I could do that. thank you again. I understand now.

OpenStudy (asnaseer):

great - I'm glad I was able to help :)